# Diffusion Autoencoders for Few-shot Image Generation in Hyperbolic Space

### v1

@@ Abstract
Few-shot image generation aims to generate diverse and high-quality images for an unseen class given only a few examples in that class. However, existing methods often suffer from a trade-off between image quality and diversity while offering limited control over the attributes of newly generated images. In this work, we propose Hyperbolic Diffusion Autoencoders (HypDAE), a novel approach that operates in hyperbolic space to capture hierarchical relationships among images and texts from seen categories. By leveraging pre-trained foundation models, HypDAE generates diverse new images for unseen categories with exceptional quality by varying semantic codes or guided by textual instructions. Most importantly, the hyperbolic representation introduces an additional degree of control over semantic diversity through the adjustment of radii within the hyperbolic disk. Extensive experiments and visualizations demonstrate that HypDAE significantly outperforms prior methods by achieving a superior balance between quality and diversity with limited data and offers a highly controllable and interpretable generation process.
@@ 1 Introduction
The advancement of generative models [62, 28, 58, 57] has succeeded in generating high-fidelity and realistic images, partially thanks to a large volume of high-quality data for model training. However, considering the widespread presence of long-tail distributions and data imbalances across image categories [31], there are many scenarios in the real world where it is impossible to obtain sufficient samples of certain categories for model training. To adapt to such practical applications, a variety of generation methods have been introduced as data augmentation techniques to generate new realistic and diverse images for a category given only a few images. This challenging task is referred to as few-shot image generation [9, 29, 31, 30, 33, 32, 15, 41, 16]. A variety of tasks can benefit from improvements in few-shot image generation, for instance, low-data detection [19] and few-shot classification [63, 68].
@@
Existing GAN-based few-shot image generation methods generally fall into three categories. Transfer-based methods [9, 43] leverage meta-learning or domain adaptation on GANs to produce new images by facilitating knowledge transfer across categories. Fusion-based methods [2, 23, 31, 30, 71, 76] combine features from multiple input images in feature space, decoding the fused features to generate images, although outputs tend to closely resemble the source images. Transformation-based methods [1, 33, 32, 15, 16] apply intra-category transformations or introduce random perturbations to samples from unseen categories, allowing image generation without extensive fine-tuning. However, these methods represent the images in the Euclidean feature space, and tend to be over-complicated, leading to generated images that frequently suffer from collapse due to limited diversity.
In contrast to previous methods that did not consider hierarchy among image categories, recent studies emphasize the importance of hierarchical structures for encoding complex visual and semantic relationships in few-shot image generation. Inspired by HAE [41], we recognize that hierarchy is prevalent in both language [49, 66, 14] and images [37, 10]. For instance, ImageNet classes [12] are based on the hierarchical structure of WordNet [18], enabling semantic hierarchies from language to be instantiated in visual data. A general prompt like “a face” corresponds to any image with a human face, while a more specific description like “a woman with sunglasses” narrows the possibilities to those with particular attributes. As shown in Fig. 1, an image can be viewed as a composition of hierarchical attributes. High-level, identity-relevant attributes define the primary identity (e.g., adding gender features to a neutral face to create “man” or “woman”), while low-level, identity-irrelevant attributes such as expressions or postures introduce variations within the category. Thus, an image can be seen as a descendant of another in the same category by adding identity-irrelevant attributes to its parent. For high-quality image generation, it is essential to capture the hierarchical structure within a large image dataset and represent it effectively in latent space. Our goal is to build a hierarchical visual representation where identity changes occur by shifting the latent code in an identity-relevant direction, and few-shot generation is achieved by moving the code in an identity-irrelevant direction.
However, existing GAN-based methods that rely on Euclidean space and its distance metrics face challenges in capturing hierarchical attribute representations, often necessitating complex disentangling mechanisms to enhance generation quality. In contrast, inspired by the application of hyperbolic geometry to images [37] and videos [64], we leverage the inherent advantages of hyperbolic space for encoding hierarchical structures. Hyperbolic space, unlike Euclidean space, serves as a continuous analog of a tree, enabling tree-like graphs to be embedded with minimal distortion [49]. This unique property provides hyperbolic space with the capacity to represent infinite semantic levels, facilitating robust and diverse attribute editing. As a result, hyperbolic space supports effective and diverse image generation for unseen categories with only a few examples, achieved through straightforward and efficient operations.
Although prior work [41] demonstrates the advantages of hierarchical representation over Euclidean latent space for diverse few-shot image generation, it faces three primary challenges (shown in Fig. 8): 1) Limited image quality due to the generative constraints of GANs, especially with insufficient training data; 2) Restricted diversity from the 1-to-1 mapping between hyperbolic and image spaces, resulting in loss of high-frequency details when latent codes are under-trained; and 3) Inability to specify the direction for editing identity-irrelevant features, often causing entanglement with identity-relevant attributes and potential category shifts in generated images. However, the recent advances in diffusion models [55, 46, 59, 73, 34, 8, 42, 72, 5] now enable high-quality image generation with rich, diverse details. Furthermore, robust pre-trained “foundation” models [58, 56] support fine-tuning with limited data and exhibit exceptional out-of-distribution generation capabilities.
@@
Based on the above findings, we propose Hyperbolic Diffusion Autoencoders (HypDAE) for few-shot image generation and editing, leveraging the natural suitability of hyperbolic space for hierarchical latent code manipulation. Our method begins by training an image encoder to capture high-level semantics, while utilizing a pre-trained Stable Diffusion (SD) model for decoding and modeling stochastic variations. A hyperbolic feature encoder and decoder are then trained to map latent vectors from Euclidean space $\mathbb{R}^{n}$ to hyperbolic space $\mathbb{D}^{n}$, with a supervised classification loss ensuring hierarchical image embeddings. Subsequently, we train a hyperbolic delta mapper [47] to project pre-trained CLIP text-image features into hyperbolic space, eliminating the need for paired text-image data. By capturing the joint attribute hierarchy between texts and images, HypDAE generates new images through two methods: 1) varying stochastic subcodes or shifting latent codes in an identity-irrelevant direction to modify category-irrelevant features, and 2) text-guided image editing along meaningful directions. Hyperbolic space enables control over the semantic diversity of generated images by adjusting the radius in the Poincaré disk, facilitating hierarchical attribute editing for flexible, high-quality, and diverse few-shot image generation and editing.
Our contributions can be summarized as follows:
@@
  • We introduce HypDAE, an innovative method for hierarchical few-shot image generation, which learns a joint hierarchical representation of text and image in hyperbolic space. To the best of our knowledge, HypDAE is the first to combine hyperbolic space and diffusion models for few-shot image generation.
@@
  • We demonstrate that in our hyperbolic latent space, semantic hierarchical relationships among images are reflected by their distances to the center of the Poincaré disk. This allows for easy interpretation of latent space and facilitates text-guided image editing.
@@
  • By leveraging pre-trained foundation models, HypDAE generates high-quality images with rich, diverse, category-irrelevant features from a single image. Extensive experiments show that HypDAE significantly outperforms existing few-shot image generation methods, offering state-of-the-art quality, diversity, and flexibility. Unlike other methods, HypDAE supports both random and customized few-shot generation and enables hierarchical image generation with varying semantic similarity.
@@ 2 Related Work
Figure 2: The overview of HypDAE. The hyperbolic autoencoder consists of a “semantic” encoder that maps the reference image to the semantic subcode ($x_{0}\rightarrow\boldsymbol{c}$), and a stable diffusion model that acts both as a “stochastic” encoder ($x_{0}\rightarrow x_{T}$) and a diffusion decoder ($(\boldsymbol{c}^{\prime},x_{T})\rightarrow x_{0}$). Here, $\boldsymbol{c}$ captures the high-level semantics while $x_{T}$ captures low-level stochastic variations, and together they can be decoded back to the original image with high fidelity. The “Hyp Encoder” is used to project the latent code from Euclidean space $\mathbb{R}^{n}$ to hyperbolic space $\mathbb{D}^{n}$ [6]. Finally, the “Hyperbolic Delta Mapper” is trained to predict meaningful editing direction in the hyperbolic latent space given text prompts. The edited semantic code is used as the condition for the diffusion decoder to generate diverse new images with the same category. For better clarity and readability, the images are shown in the pixel space $x_{T}$ instead of the VAE latent space $z_{T}$.
@@
Few-shot Image Generation. Recent advances in few-shot image generation encompass various approaches. Transfer-based methods [9, 43, 75], utilizing meta-learning or domain adaptation on GANs, often struggle with generating realistic images. Fusion-based methods, aligning random vectors with conditional images [30] or framing generation as a conditional task [23, 31, 71], generally produce outputs with limited diversity. Transformation-based approaches [1, 33, 32, 15, 16, 41], which transform single conditional images, can lack consistency. Ding et al. [15, 16] explored intra-category transformations as identity-irrelevant edits using a single sample. More recently, subject-driven generation [17, 11, 69, 24] adapts models to incorporate new concepts with limited data, exemplified by DreamBooth [59], which assigns unique identifiers to subjects using 3-5 images. This has inspired methods like Custom Diffusion [39] and further improvements [25, 7, 60]. Diffusion models (DMs) have surpassed GANs [21] in sample quality and training control but generally face slower inference and reduced flexibility due to less compact, disentangled latent spaces [61].
@@
Hyperbolic Representation Learning. Recent advancements have introduced hyperbolic space into deep learning [49, 50, 66, 64, 37], initially in NLP for hierarchical language modeling [49, 50, 66]. Riemannian optimization techniques have enabled model training within hyperbolic spaces [4, 3]. Building on this, Ganea et al. [20] developed hyperbolic adaptations of core neural network tools such as multinomial logistic regression, feed-forward, and recurrent networks. Applications of hyperbolic geometry have since expanded to image [37, 41], video [64], graph data [6, 52], and 3D shape generation [40]. Desai et al. [13] introduced a hyperbolic contrastive model for hierarchical text-image pairing, while HAE [41] demonstrated hyperbolic space’s advantages in few-shot image generation, although GAN-based results remained limited in quality. To our knowledge, this work is the first to integrate hyperbolic representation with diffusion models for high-quality few-shot image generation.
@@ 3.1 Preliminary
Hierarchical Learning in Hyperbolic Space. Our main challenge is to derive a hierarchical representation from real images that enables multi-level semantic editing, as illustrated in Fig. 1. To this end, we adopt hyperbolic space as the latent space due to its suitability for hierarchical structures. Unlike Euclidean spaces with zero curvature or spherical spaces with positive curvature, hyperbolic spaces have negative curvature, ideal for representing hierarchical data [49, 50]. Hyperbolic space, viewed as a continuous analog of trees [49], features exponential radius growth, which supports hierarchical modeling for text, images, and videos [22].
The $n$-dimensional hyperbolic space, $\mathbb{H}^{n}$, is a homogeneous, simply connected Riemannian manifold with constant negative sectional curvature [^1] We set the curvature $c=-1$.. We use the Poincaré disk model, $\left(\mathbb{D}^{n},g^{\mathbb{D}}\right)$, a preferred choice in gradient-based learning [49, 20, 50, 66, 64, 37], defined by $\mathbb{D}^{n}=\left\{x\in\mathbb{R}^{n}:\|x\|<1\right\}$ with the Riemannian metric:
@@
Identifying its tangent space $T_{x}\mathbb{D}^{n}$ with $\mathbb{D}^{n}$ allows us to transform data between Euclidean and hyperbolic spaces through the exponential and logarithmic maps with origin at $x$, respectively denoted by $\exp_{x}$ and $\log_{x}$. Explicit expressions are explained in the supplementary material.
3.2 Network Architecture
The overall pipeline of HypDAE, shown in Fig. 2, is divided into three stages for two key reasons: 1) Multiple loss functions must be optimized for various image generation functionalities, and joint end-to-end optimization is challenging for satisfactory results; 2) End-to-end training requires extensive labeled data, which is difficult to obtain. By staging the pipeline, we can leverage existing datasets collected for general vision tasks and avoid the need for new task-specific paired data. Specifically: 2) In stage II, we train a hyperbolic encoder and decoder to map visual features from stage I into hyperbolic space using supervised classification loss, then reconstruct these features back to Euclidean space. This step captures the hierarchy across the image corpus, enabling the generation of new images with varied identity-irrelevant features by moving latent codes within the category. 3) In stage III, a hyperbolic delta mapper aligns CLIP text features with image features from stage II within the hyperbolic space, facilitating interpretable direction translation in latent space and disentangled, text-guided image editing for controllable few-shot image generation.

### v2

# HypDAE: Hyperbolic Diffusion Autoencoders for Hierarchical Few-shot Image Generation

@@ Abstract
Few-shot image generation aims to generate diverse and high-quality images for an unseen class given only a few examples in that class. A key challenge in this task is balancing category consistency and image diversity, which often compete with each other. Moreover, existing methods offer limited control over the attributes of newly generated images. In this work, we propose Hyperbolic Diffusion Autoencoders (HypDAE), a novel approach that operates in hyperbolic space to capture hierarchical relationships among images from seen categories. By leveraging pre-trained foundation models, HypDAE generates diverse new images for unseen categories with exceptional quality by varying stochastic subcodes or semantic codes. Most importantly, the hyperbolic representation introduces an additional degree of control over semantic diversity through the adjustment of radii within the hyperbolic disk. Extensive experiments and visualizations demonstrate that HypDAE significantly outperforms prior methods by achieving a better balance between preserving category-relevant features and promoting image diversity with limited data. Furthermore, HypDAE offers a highly controllable and interpretable generation process. Code is available at: https://github.com/lingxiao-li/HypDAE.
@@ 1 Introduction
Generative models [54, 24, 50, 49] have succeeded in generating high-fidelity and realistic images, partially thanks to a large volume of high-quality data for model training. However, with the widespread presence of long-tail distributions and data imbalances across image categories [27], there are many scenarios in the real world where it is impossible to collect sufficient samples of certain categories for model training. It is difficult for generative models trained on well-sampled categories to generate realistic and diverse images for a novel category given only a few examples. This challenging task is known as few-shot image generation, which aims to synthesize images that preserve the category-level identity of the limited input samples [8, 25, 27, 26, 29, 28, 13, 37, 14].
@@
Existing few-shot image generation methods are primarily GAN-based and fall into three categories: transfer-based approaches [8, 38], which use meta-learning or domain adaptation for cross-category generalization but often face limited transferability; fusion-based approaches [2, 19, 27, 26, 61, 66], which fuse features from multiple exemplars but tend to produce outputs overly similar to the inputs; and transformation-based approaches [1, 29, 28, 13, 14], which apply intra-category perturbations without fine-tuning but often lack diversity. Recent diffusion-based methods [33, 51, 5] for object-level personalization focus on instance identity, require test-time fine-tuning, and depend on prompt engineering, making them incompatible with our category-level setting without textual inputs or model updates. However, all these methods operate in Euclidean feature space, which limits their ability to capture hierarchical semantic structure—a key requirement for generalizing from few examples in category-level few-shot image generation.
In contrast to prior methods, recent work such as HAE [37] highlights the importance of modeling hierarchical structures in few-shot image generation. Similar to language [41, 57, 12], images also exhibit semantic hierarchies [31, 9], where each image can be viewed as a composition of attributes at different levels. As illustrated in Fig.˜1, high-level, identity-relevant attributes (e.g., gender or age) define the semantic core of a category, while low-level, identity-irrelevant attributes (e.g., expression or hairstyle) introduce intra-class variation. Capturing this hierarchical organization is essential for generating diverse yet category-consistent images. Hyperbolic space provides a natural embedding for such structures, as it can represent tree-like relationships with low distortion [41]. This enables infinite layers of semantics to be encoded compactly, allowing identity-preserving edits along radial directions and intra-category diversity through tangential shifts in the latent space, which facilitates a structured and interpretable latent space that supports controllable few-shot image generation.
Despite the advantages of hierarchical representation for few-shot image generations [37], existing methods are dominated by GAN-based approaches and face three primary challenges: 1) Suboptimal image quality due to the generative constraints of GANs, particularly with insufficient training data; 2) Reduced diversity from the 1-to-1 mapping between hyperbolic and image spaces, which leads to the loss of high-frequency details when latent codes are insufficiently trained; and 3) The necessity of labeled data for learning hierarchical latent representations. In parallel, recent advances in diffusion models [47, 33, 51, 62, 5] enable high-quality image generation with rich, diverse details. Moreover, pre-trained foundation models (e.g., Stable Diffusion [50], CLIP [48]) further support adaptation with limited data and strong generalization.
@@
To address the challenges, we propose Hyperbolic Diffusion AutoEncoders (HypDAE), a novel method for leveraging the natural suitability of hyperbolic space for hierarchical latent code manipulation and the generative ability of diffusion models for few-shot image generation. In particular, we separate the representation of given images into two subcodes [47]. Our method begins by training an image encoder to capture high-level semantic subcode, while utilizing a pre-trained Stable Diffusion (SD) [50] model for decoding and modeling stochastic subcode by reversing the generative process. A hyperbolic feature encoder and decoder are then trained to map latent vectors from Euclidean space $\mathbb{R}^{n}$ to hyperbolic space $\mathbb{D}^{n}$, with classification loss ensuring hierarchical image embeddings. Finally, to reduce the data labeling cost for training a hyperbolic encoder, pseudo-labeling is enabled by zero-shot classification with pre-trained vision models. By capturing the attribute hierarchy among images, HypDAE generates new images through two methods: 1) varying stochastic subcode of the Diffusion Autoencoders, and 2) randomly shifting semantic subcode in an identity-irrelevant direction to modify category-irrelevant features. Hyperbolic space enables control over the semantic diversity of generated images by adjusting the radii in hyperbolic space. This facilitates hierarchical attribute editing for flexible, high-quality, and diverse few-shot image generation.
Our contributions are summarized as follows:
@@
  • We introduce HypDAE, an innovative method for few-shot image generation, which learns the hierarchical representation of images in hyperbolic space. To the best of our knowledge, HypDAE is the first to enable diffusion models to model semantic hierarchical information in hyperbolic space for few-shot image generation.
@@
  • We demonstrate that in our hyperbolic latent space, semantic hierarchical relationships among images are reflected by their distances to the center of the Poincaré disk. This allows for easy interpretation of the latent space.
@@
  • By leveraging pre-trained foundation models, HypDAE generates high-quality images with rich, diverse, category-irrelevant features from a single image. Extensive experiments show that HypDAE significantly outperforms existing few-shot image generation methods without requiring human-annotated class labels, offering state-of-the-art quality, diversity, and flexibility. This demonstrates the robustness of HypDAE under settings with automatically generated labels, making it a strong solution for label-free few-shot image generation.
@@ 2 Related Work
Few-shot Image Generation. Given a few samples from one novel category, few-shot image generation aims to produce more new images with high quality and various diversity. Recent advances in few-shot image generation encompass various approaches. Transfer-based methods [8, 38, 65], utilizing meta-learning or domain adaptation on GANs, often struggle with generating realistic images. Fusion-based methods, aligning random vectors with conditional images [26] or framing generation as a conditional task [19, 27, 61], generally produce outputs with limited diversity. Transformation-based approaches [29, 28, 37], which transform single conditional images, can lack consistency. Ding et al. [13, 14] explored intra-category transformations as identity-irrelevant edits using a single sample. More recently, Diffusion-based subject-driven generation [15, 10, 59, 20] adapts models to incorporate new concepts with limited data, exemplified by DreamBooth [51], which assigns unique identifiers to subjects using 3-5 images. This has inspired methods like Custom Diffusion [33] and further improvements [21, 7, 52]. However, these works focus on instance-level identity preservation, require test-time fine-tuning, and rely heavily on prompt engineering. Therefore, their objectives, problem settings, and evaluation protocols differ fundamentally from our task.
@@
Hyperbolic Representation Learning. Recent advancements have introduced hyperbolic space into deep learning [41, 42, 57, 55, 31], initially in NLP for hierarchical language modeling [41, 42, 57]. Riemannian optimization techniques have enabled model training within hyperbolic spaces [4, 3]. Building on this, Ganea et al[16] developed hyperbolic adaptations of core neural network tools such as multinomial logistic regression, feed-forward, and recurrent networks. Applications of hyperbolic geometry have since expanded to image [31, 37], video [55], graph data [6, 44], and 3D shape generation [36]. Desai et al[11] introduced a hyperbolic contrastive model for hierarchical text-image pairing, while HAE [37] demonstrated hyperbolic space’s advantages in few-shot image generation, although GAN-based results remained limited in quality. To our knowledge, this work is the first to integrate hyperbolic representation with diffusion models for high-quality few-shot image generation.
Figure 2: Illustration of the property of hyperbolic space on the Poincaré disk. Given two latent codes of Maltese dog $\boldsymbol{c}_{h}^{\prime}$ and $\boldsymbol{c}_{h}^{\prime\prime}$ on the edge of Poincaré disk, the geodesic between these two points is the red curve rather than a straight line in Euclidean space. Therefore, their average latent code is calculated as $\bar{\boldsymbol{c}_{h}}$, which is closer to the center. Thus, $\bar{\boldsymbol{c}_{h}}$ can be viewed as the “parent" of $\boldsymbol{c}_{h}$, $\boldsymbol{c}_{h}^{\prime}$, and $\boldsymbol{c}_{h}^{\prime\prime}$. One can generate diverse images without changing the category by moving the latent code from one child to another of the same parent in the hyperbolic space.
Figure 3: The overview of HypDAE. The hyperbolic autoencoder consists of a “semantic” encoder that maps the reference image to the semantic code ($x\rightarrow\boldsymbol{c}$), and a stable diffusion model that acts both as a “stochastic” encoder ($x\rightarrow\boldsymbol{z}_{T}$) and a diffusion decoder ($(\boldsymbol{c}^{\prime},\boldsymbol{z}_{T})\rightarrow x^{\prime}$). Here, $\boldsymbol{c}$ captures the high-level semantics while $\boldsymbol{z}_{T}$ captures low-level stochastic variations; they can be decoded back with high fidelity. The “Hyp Encoder” is used to project the latent code from Euclidean space $\mathbb{R}^{n}$ to hyperbolic space $\mathbb{D}^{n}$ [6]. The semantic code can be edited as Fig.˜2 shows and be used as the condition for the diffusion decoder to generate diverse new images with the same category. The VAE module of SD is omitted for better illustration.
@@ 3 Method
To generate diverse new images from a few reference images while preserving their identity, it is essential for our model to capture both identity-relevant features (to maintain identity) and identity-irrelevant features (to enhance diversity). To achieve this, we design a diffusion autoencoder that extracts elementary identity-relevant features via a semantic encoder and identity-irrelevant features via a stochastic encoder. Furthermore, a hyperbolic encoder-decoder is introduced to further disentangle high-order semantic features. Unlike HAE, which applies hyperbolic representations in GANs, our method integrates hyperbolic semantics into diffusion models. We introduce a stable training pipeline, and encode hyperbolic embeddings as diffusion conditions, enabling scalable and interpretable few-shot generation.
We give details of our method in the following sections. Specifically, Sec.˜3.1 details the modeling of hierarchical representations, Sec.˜3.2 elaborates on the framework and loss functions of HypDAE, and Sec.˜3.3 and Sec.˜3.4 discuss the pseudo-labeling process and hyperbolic latent editing algorithms, respectively.
@@ 3.1 Preliminary
Hierarchical Learning in Hyperbolic Space. A key challenge in multi-level semantic editing is deriving a hierarchical representation from real images, as illustrated in Fig.˜2. To address this, we utilize hyperbolic space as the latent space, given its suitability for hierarchical structures. Unlike Euclidean spaces (zero curvature) or spherical spaces (positive curvature), hyperbolic spaces exhibit negative curvature, making them ideal for modeling hierarchical data [41, 42]. As a continuous analog of trees [41], hyperbolic space enables hierarchical representation through its exponential radius growth, facilitating structured modeling across text, images, and videos [18].
The $n$-dimensional hyperbolic space, $\mathbb{H}^{n}$, is a homogeneous, simply connected Riemannian manifold with constant negative sectional curvature [^1] We set the curvature c=1c=-1 in this paper.. We use the Poincaré disk model following [31], $\left(\mathbb{D}^{n},g^{\mathbb{D}}\right)$, a preferred choice in gradient-based learning [41, 16, 42, 57, 55, 31], defined by $\mathbb{D}^{n}=\left\{x\in\mathbb{R}^{n}:\|x\|<1\right\}$ with the Riemannian metric:
@@
A geodesic is defined as a locally minimized-length curve between two points, as illustrated in Fig.˜2. The midpoint of this geodesic, closer to the origin, represents the mean of two latent codes in hyperbolic space. This property ensures that the mean of two leaf embeddings is not another leaf but their hierarchical parent, as illustrated in Fig.˜2 on a 2-D Poincaré disk [55]. Embeddings near the disk’s edge correspond to fine-grained images, whereas those closer to the center represent more abstract features, such as an average face.
3.2 Hyperbolic Diffusion AutoEncoders (HypDAE)
The overall pipeline of HypDAE, illustrated in Fig.˜3, consists of two stages: 1) In Stage I, a semantic encoder maps input images to meaningful latent codes, serving as conditions for the pre-trained Stable Diffusion (SD). To enhance diversity and prevent direct replication, a content bottleneck and strong augmentation techniques are applied. 2) In Stage II, a hyperbolic encoder-decoder projects visual features from Stage I into hyperbolic space using supervised classification loss and then reconstructs them in Euclidean space. This process captures hierarchical relationships within the image corpus, facilitating the generation of diverse images by manipulating latent codes within the same category.
We design a two-stage model for two reasons: 1) jointly and end-to-end optimizing the whole pipeline is challenging for satisfactory results as multiple loss functions must be optimized for various image generation functionalities; 2) end-to-end training requires extensive labeled data, which is difficult to obtain. By staging the pipeline, we can leverage existing datasets collected for general vision tasks and avoid the need for new task-specific paired data.
Diffusion Autoencoders. In this stage, we aim to generate new images $x^{\prime}$ of the reference image $x$ with rich stochastic details (e.g., hair, fur, color, etc.) while maintaining the identity and category of the image unchanged.